Question  9

If the production function is given by


Q = 300L − 4L,


where Q denotes output and L denotes the size of workforce, calculate the value of marginal product of  labour if L = 9.

a.        11

b.        46

c.        16

d.    146


1
Expert's answer
2021-06-14T18:56:44-0400

Solution.

"Q=300\sqrt{l}-4l."

Marginal product

"MP=\frac{dQ}{dl}."

"MP=300\frac{1}{2\sqrt{l}}-4=\frac{150}{\sqrt{l}}-4."

If "l=9," then

"MP=\frac{150}{\sqrt{9}}-4=46."

Answer. b